# Exploring the Naturalness of AI-Generated Images

### 14 Dec 2023 v2

@@ Abstract
The proliferation of Artificial Intelligence-Generated Images (AGIs) has greatly expanded the Image Naturalness Assessment (INA) problem. Different from early definitions that mainly focus on tone-mapped images with limited distortions (e.g., exposure, contrast, and color reproduction), INA on AI-generated images is especially challenging as it has more diverse contents and could be affected by factors from multiple perspectives, including low-level technical distortions and high-level rationality distortions. In this paper, we take the first step to benchmark and assess the visual naturalness of AI-generated images. First, we construct the AI-Generated Image Naturalness (AGIN) database by conducting a large-scale subjective study to collect human opinions on the overall naturalness as well as perceptions from technical and rationality perspectives. AGIN verifies that naturalness is universally and disparately affected by both technical and rationality distortions. Second, we propose the Joint Objective Image Naturalness evaluaTor (JOINT), to automatically learn the naturalness of AGIs that aligns human ratings. Specifically, JOINT imitates human reasoning in naturalness evaluation by jointly learning both technical and rationality perspectives. Experimental results show our proposed JOINT significantly surpasses baselines for providing more subjectively consistent results on naturalness assessment. Our database and code will be released in https://github.com/zijianchen98/AGIN.
@@ 1 Introduction
Figure 1: Exemplar images with diverged technical quality and rationality. Evaluating image naturalness from different perspectives (technical/rationality) may produces biases, motivating us to construct AGIN, the first INA database with opinions from multi-perspectives.
@@
Conventionally, image naturalness is described as the degree of correspondence between a real-life scene and a photograph displayed on a device based on some technical criteria (e.g., texture, exposure, color reproduction, shooting artifacts) [18, 8, 14], which has been utilized for image quality assessment (IQA) to compare and guide the optimization of systems and algorithms [24, 93, 44]. Under this perspective, the image with richer details (Fig. 1c, Fig. 1d) should have notably better naturalness than blur image in Fig. 1a. More recently, the emergence of AGI broadens the definition of image naturalness to comprise more non-technical semantic factors (e.g., existence, layout), which is normally regarded as the rationality perspective [59, 45]. However, it is highly subjective and its mechanism of how rationality affects human perception in image naturalness reasoning, i.e., human-naturalness opinions, is still ambiguous and may be multidimensionally coupled.
@@
In this paper, we make the first attempt to evaluate naturalness of AI-generated images, a new field of quality assessment with increasing attention [68, 50, 91, 109, 12]. To benchmark naturalness of AI-generated images, we contribute AI-Generated Image Naturalness (AGIN) database, the first-of-its-kind database to study this problem. Specifically, AGIN contains 6,049 images collected from five different generative tasks with 18 model variants to ensure diversity. A total of 907,350 human opinions for technical and rationality perspectives as well as their effects on overall naturalness scores were collected from 30 participants. Since the perspectives and factors studied in our research are related to common IQA problems and not limited to AI-generated images, our methods and insights also can be applicable to other forms of multimedia.
@@
AGIN provides several valuable observations for understanding human reasoning in visual naturalness. Firstly, we find both technical distortions (e.g., contrast, blur, and generative artifacts) and rationality distortions (e.g., existence, color, and layout) can affect visual naturalness significantly. The proportion of technical and rationality factors for AGIs with worse naturalness scores ($\mathrm{MOS}\in[1,3]$) is about 1:1.17. Secondly, we notice that most factors in the two perspectives are relevant, but have disparate impacts on the naturalness score, which can result in biased naturalness assessment. Furthermore, we also observe that the overall naturalness score can be well-approximated by a linear weighted sum of technical score and rationality score ($\mathrm{MOS}=0.145\mathrm{MOS_{T}}+0.769\mathrm{MOS_{R}}$). This correlation suggests joint learning of technical and rationality branches can be a feasible way to model naturalness.
@@
Based on AGIN database, we propose the Joint Objective Image Naturalness evaluaTor (JOINT), an objective naturalness assessment method that offers high alignment with human perception. JOINT aims to mimic human reasoning of image naturalness by jointly learning on both technical and rationality branches. Specifically, given the different characteristics of each branch, we elaborate several designs, such as patch partition, deep feature regularization, and pretraining, to allocate each branch with corresponding learning interests. Two different supervision schemes including using the overall naturalness scores (JOINT) and the respective scores for each perspective (JOINT++), are applied to train the model. Finally, we use an effective subjective weighting strategy combined with the predictions of two branches to compute the overall naturalness score. Experimental results on AGIN database verify the effectiveness of our proposed JOINT and JOINT++ that not only outperform baselines on the overall naturalness assessment but also provide more subjectively consistent results for technical and rationality perspectives.
@@ 2 Related Work
AI-generated Image and Naturalness. Generative models have emerged as an effective paradigm for image synthesis [74, 76, 66, 73, 62, 101]. Nonetheless, most generative adversarial network (GAN)-based methods [104, 70, 99] are prone to produce visually unnatural results due to their instability and mode collapse issues. Even state-of-the-art diffusion-based generative models [41, 92, 60, 6] introduce oftentimes perceptible unnatural perturbations such as spurious details, disordered layout, and color mismatch on images. Prior naturalness prediction approaches [24, 55, 93, 25, 111], driven by image statistic distribution, have predominantly focused on natural scene images (NSIs), which fail in AGIs, where exist diverse contextual content variations with less significant intrinsic properties (e.g. resolution, color space, and image format). As a result, it is challenging to design an effective image naturalness assessment method for AI-generated images that can be used to optimize the naturalness of the generated images and make them more robust to real-world applications.
@@
AI-generated Image Assessment. Existing AI-generated image assessment research mainly focuses on the perceptual quality. Early objective metric such as Inception Score (IS) [77] measures perceptual quality by calculating the uniformity of AGIs group features from the output of Inception model. Distance-based methods such as Fréchet Inception Distance (FID) [31] and Kernel Inception Distance (KID) [4] as well as Precision-Recall [42] evaluate the discrepancy between distributions of AGI and NSI. Nevertheless, the above methods are all group-targeted and not suitable for assessing single image. Besides, the widely used CLIPScore [30] is already saturated in comparing state-of-the-art generative models with authentic images and can inflate for a model trained to optimize text-to-image alignment in the CLIP space [68]. This empirical evidence of the failure of the automatic measures motivates human evaluation of perceived quality. Kirstain et al. [40] collect human preferences between two generated images from the same prompt for text-to-image tasks. Wang et al. [85] investigate the impact of quality, fidelity, and correspondence of AI-generated images on human visual perception. Similarly, Li et al. [47] conduct a subjective evaluation to annotate images from both perception and alignment dimensions with varying input prompts and internal parameters in AGI models. However, these studies, which merely collect coarse, single-voice, and overall subjective opinions, lack the exploration of specific factors with fine-grained and explainable evaluations on various generation tasks.
Figure 2: Pie chart of AGIN database content distribution. The blue, green, red, orange, purple, and black partitions indicate that the images are derived from text-to-image, image translation, image inpainting, image colorization, image editing, and real-world, respectively.
@@ 3 AI-Generated Image Naturalness Database
In this section, we elaborate on the construction procedures of the proposed AGIN database, along with the subjective human evaluation (Fig. 3). The database includes 6,049 AI-generated images, on which we collected 907,350 ratings in terms of the overall naturalness score and its two perspectives: the technical and rationality score, as well as their respective main influencing factor.
@@ 3.1 Data Preparation
As an initial investigation, we choose five sources of AI-generated images including text-to-image, image translation, image inpainting, image colorization, and image editing, that typically suffer from naturalness problems. We select 18 models including: (1) text-to-image, 5 models with over 400 prompts are used for image generation. (2) image translation, 5 models with various text-, image-, or mask-guided (e.g. edge map, semantic map) translation. (3) image inpainting, 2 models that take mutilated images as inputs. (4) image colorization, 3 models that colorize the grayscale images. (5) image editing, 3 models that perform layout or content editing on the image via text prompts or interactive anchor points. In addition to the AI-generated images, we also added 200 extra real images into the AGIN database to help analyze the accuracy of the subjective experiment and objective algorithms, which stand for a high level of naturalness. The distribution of the selected models and categories is displayed in the pie chart of Fig. 2. We carefully reproduce the generation processes in supplementary materials.
@@ 3.2 Design of the Human Evaluation
3.2.1 Choice of Naturalness-related Factors
We conduct the evaluation from two perspectives, i.e., the low-level technical perspective and the high-level rationality perspective, as follows.
@@ 3.2.1 Choice of Naturalness-related Factors
(T-1) Luminance: Unrecognizable regions of images due to extremely high/low brightness [89].
@@
Figure 3: Workflow of the in-lab human evaluation in AGIN. It mainly includes four stages: 1) In-Lab Training, 2) Instruction, 3) Annotation, and 4) Testing, discussed in Sec. 3.2.2.
3.2.2 Design of the Study
Participants and Apparatus. To ensure the comprehensiveness and reliability of the evaluation, we recruited 30 participants (18 male, 12 female, age=22.6±3.1) from campus, all with normal (corrected) eyesight. We conduct the subjective studies in-lab to ensure that all subjects have a clear and consistent understanding of all factors. Each participant is compensated \$240 for evaluating 6,049 images. All images are displayed on a 27-inch screen with a resolution of 2560$\times$1440 and a viewing distance of about 70cm.
Rating Strategy and Wording. We discuss the concrete form for human evaluation as follows. 1) Task-oriented absolute choice. Since the wording of questions and labels can significantly affect annotators’ labeling behavior, we abandon the traditional 3-point or 5-point Likert scale that only provides endpoint labels from worst to best and is too vague to describe the degree of naturalness. As a solution, we elaborate the questions and labels for three different perspectives to reduce subjectivity rather than using general ones (e.g. bad, good, or very good). 2) Pick up the main factor. Many existing subjective studies merely focus on the assessment of the overall score but neglect to explore the underlying factors. Therefore, we ask subjects to choose a primary factor that affects most for each perspective after rating the general scores, which enables us to investigate the correlation between each dimension and image naturalness.
Training, Testing, and Annotation. The workflow of the human evaluation is illustrated in Fig. 3. Before conducting the formal study, we manually generated 10 exemplars beyond the AGIN database for each dimension as training images to familiarize subjects with the goal of this evaluation. Subsequently, we instruct the subjects to rate the technical quality, rationality, and overall naturalness of each image from $\left\{1,2,3,4,5\right\}$, and select the main factors that most affect the technical quality and rationality perspectives. For testing, we randomly insert 10 golden images into each session as an inspection to ensure the quality of annotation. We defer more details of human evaluation and quality control of ratings to supplementary materials.
Figure 4: Data properties of AGIN. (a) The correlations between technical and rationality perspectives, and (b) distributions of overall naturalness scores.
Figure 5: Proportions for each dimension. The tendency of main factors chosen by participants across different ranges of naturalness scores.
@@ 3.3 Insights
We first provide a visualization of the data properties in AGIN database (Fig. 4a), from which we can observe the correlation between the technical perspective and the rationality perspective. Tab. 1 lists the quantitative results of Spearman and Pearson correlation between different perspectives, where the mean technical score, mean rationality score, and mean naturalness score are denoted as $\mathrm{MOS_{T}}$, $\mathrm{MOS_{R}}$, and $\mathrm{MOS}$, respectively. It can be observed that the two perspectives affect naturalness unequally (i.e., rationality has a greater impact on the overall naturalness than technical perspective, while the linear weighted sum of two perspectives can better approximate the overall naturalness than any single form). This could lead to biased naturalness evaluation unwittingly when using mainstream IQA models that follow the overall MOS regression strategy.
Insight ❷: Factors in two perspectives are related, but have disparate impacts on naturalness scores.
Fig. 5 shows the proportion of each factor in different ranges of naturalness scores, where T-Null and R-Null indicate that there exist no factors affecting naturalness or the subjects have difficulty in choosing the main factors. First, T-Null and R-Null are more prevalent in images with preferable naturalness ($\mathrm{MOS}\in[4,5]$), indicating the impact of technical and rationality perspectives on naturalness. Additionally, we notice that humans are more sensitive to generated artifacts (T-5) and blur (T-4) in terms of technical quality while focusing more on the existence (R-1) of the image contents in terms of rationality. Furthermore, we find especial high proportions of artifacts (T-5), existence (R-1), and layout (R-3) in the case of poor naturalness ($\mathrm{MOS}\in[1,2]$), suggesting that they are important naturalness factors for AGIs to take into account. Considering the inter-relation between technical quality and rationality, we speculate that severe artifact distortions may lead to irrational content and chaotic layout. We also provide examples with varying degrees of effect for each dimension in Fig. 6, to better illustrate the manifestations of the naturalness problem in different dimensions. Overall, these newly contributed dimensions describe naturalness concerns of AGIs, some of which have never been encountered in conventional IQA domain, providing reliable intuitions for developing objective naturalness assessment models.
@@
$\mathrm{MOS_{T}+MOS_{R}}$ $\mathrm{aMOS_{T}+bMOS_{R}}$
@@
Table 1: Correlation of perspectives. The SRCC and PLCC between different perspectives and overall naturalness (MOS) for all 6,049 images in AGIN. $a$ and $b$ are $0.145$ and $0.769$, respectively.
Figure 7: Framework of the proposed JOINT and JOINT++ with joint learning strategy. It consists of the technical prior branch (Sec. 4.1) and the rationality perceiving branch (Sec. 4.2). More details can be found in supplementary materials.

### 04 Mar 2024 v3

# https://github.com/zijianchen98/AGIN

@@ Abstract
The proliferation of Artificial Intelligence-Generated Images (AGIs) has greatly expanded the Image Naturalness Assessment (INA) problem. Different from early definitions that mainly focus on tone-mapped images with limited distortions (e.g., exposure, contrast, and color reproduction), INA on AI-generated images is especially challenging as it has more diverse contents and could be affected by factors from multiple perspectives, including low-level technical distortions and high-level rationality distortions. In this paper, we take the first step to benchmark and assess the visual naturalness of AI-generated images. First, we construct the AI-Generated Image Naturalness (AGIN) database by conducting a large-scale subjective study to collect human opinions on the overall naturalness as well as perceptions from technical and rationality perspectives. AGIN verifies that naturalness is universally and disparately affected by technical and rationality distortions. Second, we propose the Joint Objective Image Naturalness evaluaTor (JOINT), to automatically predict the naturalness of AGIs that aligns human ratings. Specifically, JOINT imitates human reasoning in naturalness evaluation by jointly learning both technical and rationality features. We demonstrate that JOINT significantly outperforms baselines for providing more subjectively consistent results on naturalness assessment.
Keywords:
AI-generated Images Image Naturalness Assessment Database
Figure 1: The proposed AGIN, first-of-this-kind image naturalness assessment database with human opinions from technical, rationality, and overall naturalness perspectives, focusing on 5 generative tasks (i.e., text-to-image, image translation, image inpainting, image colorization, and image editing).
Figure 2: The motivation of naturalness assessment for AI-generated images: multi-perspective settings can effectively avoid the perceptual bias on single absolute evaluation, and provide more accurate judgments to serve as downstream supervision.
@@ 1 Introduction
Conventionally, image naturalness is described as the degree of correspondence between a real-life scene and a photograph displayed on a device based on some technical criteria (e.g., texture, exposure, color reproduction, shooting artifacts) [75, 8, 14], which has been utilized for image quality assessment (IQA) to compare and guide the optimization of systems and algorithms [22, 94, 43]. Under this theory, the images with richer details (Fig. 2(c) and Fig. 2(d)) should have notably better naturalness than the blurred image in Fig. 2(a), which is opposite to the human opinion. More recently, the emergence of AGI broadens the definition of image naturalness to comprise more non-technical semantic factors (e.g., existence, layout), which are normally regarded as rationality perspective [58, 44]. However, it is highly subjective and its mechanism of how rationality affects human perception in image naturalness reasoning, i.e., human-naturalness opinions, is still ambiguous and may be multidimensionally coupled.
@@
In this paper, we make the first attempt to evaluate the naturalness of AI-generated images, a new field of quality assessment with increasing attention [67, 49, 92, 110, 13]. To benchmark the naturalness of AI-generated images, we contribute AI-Generated Image Naturalness (AGIN) database, the first-of-its-kind database to study this problem. Specifically, AGIN contains 6,049 images collected from five different generative tasks with 18 model variants to ensure diversity. A total of 907,350 human opinions for technical and rationality perspectives as well as their effects on overall naturalness scores were collected from 30 participants. Since the perspectives and factors studied in our research are related to common IQA problems and not limited to AI-generated images, our methods and insights also can be applicable to other forms of multimedia.
@@
AGIN provides several valuable observations for understanding human reasoning in visual naturalness. Firstly, we find both technical distortions (e.g., contrast, blur, and generative artifacts) and rationality distortions (e.g., existence, color, and layout) can affect visual naturalness significantly. The proportion of technical and rationality factors for AGIs with worse naturalness scores ($\mathrm{MOS}\in[1,3]$) is about 1:1.17. Secondly, we notice that most factors in the two perspectives are relevant, but have disparate impacts on the naturalness score, which can result in a biased naturalness assessment. Furthermore, we also observe that the overall naturalness score can be well-approximated by a linear weighted sum of technical score and rationality score ($\mathrm{MOS}=0.145\mathrm{MOS_{T}}+0.769\mathrm{MOS_{R}}$). This correlation suggests joint learning of technical and rationality branches can be a feasible way to predict naturalness.
@@
With the AGIN database, we propose the Joint Objective Image Naturalness evaluaTor (JOINT), an objective naturalness assessment method that offers high alignment with human perception. JOINT aims to mimic human reasoning of image naturalness by jointly learning on both technical and rationality branches. Specifically, given the different characteristics of each branch, we elaborate several designs, such as patch partition, deep feature regularization, and pretraining, to allocate each branch with corresponding learning interests. Two different supervision schemes including using the overall naturalness scores (JOINT) and the respective scores for each perspective (JOINT++) are applied to train the model. Finally, we use an effective subjective weighting strategy combined with the predictions of two branches to compute the overall naturalness score. Experimental results on AGIN database verify the effectiveness of our proposed JOINT and JOINT++ that not only outperform baselines on the overall naturalness assessment but also provide more subjectively consistent results for technical and rationality perspectives.
@@ 2 Related Work
AI-generated Image and Naturalness. Generative models have emerged as an effective paradigm for image synthesis [76, 78, 65, 73, 61, 103]. Nonetheless, most generative adversarial network (GAN)-based methods [105, 69, 100] are prone to produce visually unnatural results due to their instability and mode collapse issues. Even state-of-the-art diffusion-based generative models [40, 93, 59, 6] introduce oftentimes perceptible unnatural perturbations such as spurious details, disordered layout, and color mismatch on images. Prior naturalness prediction approaches [22, 54, 94, 23, 112], driven by image statistic distribution, have predominantly focused on natural scene images (NSIs), which fail in AGIs, where exist diverse contextual content variations with less significant intrinsic properties (e.g.resolution, color space, and image format). As a result, it is challenging to design an effective naturalness assessment method for AI-generated images that can be used to optimize the naturalness of the generated images and make them more robust to real-world applications.
@@
AI-generated Image Assessment. Existing AI-generated image assessment research mainly focuses on perceptual quality. Early objective metrics such as Inception Score (IS) [79] measures perceptual quality by calculating the uniformity of AGI group features from the output of Inception model. Distance-based methods such as Fréchet Inception Distance (FID) [29] and Kernel Inception Distance (KID) [4] as well as Precision-Recall [41] evaluate the discrepancy between distributions of AGI and NSI. Nevertheless, the above methods are all group-targeted and not suitable for assessing single image. Besides, the widely used CLIPScore [28] is already saturated in comparing state-of-the-art generative models with authentic images and can inflate for a model trained to optimize text-to-image alignment in the CLIP space [67]. This empirical evidence of the failure of the automatic measures motivates human evaluation of perceived quality. Kirstain et al. [39] collected human preferences between two generated images from the same prompt for text-to-image tasks. Wang et al. [87] investigated the impact of quality, fidelity, and correspondence of AI-generated images on human visual perception. Similarly, Li et al. [46] conducted a subjective evaluation to annotate images from both perception and alignment dimensions with varying input prompts and internal parameters in AGI models. However, these studies, which merely collect coarse, single-voice, and overall subjective opinions, lack the exploration of specific factors with fine-grained and explainable evaluations on various generation tasks.
Database Image Source #Content #Image Perspective Distortion
LIVE (2004) [80] Kodak test set 30 779 Quality 5 Artificial
TID2008 (2008)[71] Kodak test set 25 1,700 Quality 17 Artificial
CSIQ (2009)[42] Kodak test set 30 866 Quality 6 Authentic
TID2013 (2013)[70] Kodak test set 25 3,000 Quality 24 Artificial
LIVEC (2015)[19] Camera 1162 1162 Quality 15 Authentic
WED (2017)[60] Internet 4,744 94,880 Quality 4 Artificial
MDID (2017)[83] Internet 20 1,600 Quality 5 Artificial
PieAPP (2018)[72] WED [60] 200 20,280 Quality 75 Artificial
KADID-10k (2019)[52] Internet 81 10,125 Quality 25 Artificial
KonIQ-10k (2020)[30] Multimedia 10,073 10,073 Quality -- Authentic
PIPAL (2020)[33] DIV2K [1], Flickr2K [84] 250 29,000 Quality 40 Artificial
AGIQA-1k (2023) [110] AI-generated 1,080 1,080 Quality 2 Generative
AIGCIQA2023 (2023) [87] AI-generated 2,400 2,400 Quality, Authenticity, Correspondence 6 Generative
AGIQA-3k (2023) [46] AI-generated 2,982 2,982 Quality, Alignment 6 Generative
AGIN (Ours) AI-generated 6,049 6,049 Technical, Rationality, Naturalness 18 Generative
@@ 3 AI-Generated Image Naturalness Database
In this section, we elaborate on the construction procedures of the proposed AGIN database, along with the subjective human evaluation (Fig. 3). The database includes 6,049 AI-generated images, upon which we collected 907,350 ratings in terms of the overall naturalness score and its two perspectives: the technical and rationality scores, as well as their respective main influencing factor. A quick comparison of related datasets can be found in Tab. 1.
@@ 3.1 Data Preparation
As an initial investigation, we choose five sources of AI-generated images from text-to-image, image translation, image inpainting, image colorization, and image editing tasks, which typically suffer from naturalness problems. We select 18 models including: (1) text-to-image, 5 models with over 400 prompts are used for image generation. (2) image translation, 5 models with various text-, image-, or mask-guided (e.g. edge map, semantic map) translation. (3) image inpainting, 2 models that take mutilated images as inputs. (4) image colorization, 3 models that colorize the grayscale images. (5) image editing, 3 models that perform layout or content editing on the image via text prompts or interactive anchor points. In addition to the AI-generated images, we also added 200 extra real images into the AGIN database to help analyze the accuracy of the subjective experiment and objective algorithms, which stand for a high level of naturalness. The distribution of the selected models and categories is displayed in the pie chart of Fig. 1. We carefully reproduce the generation processes in supplementary materials.
Figure 3: Workflow of the human evaluation in AGIN: source images are first collected from 5 generative tasks and real-world image datasets (a), and then we conduct in-lab training with instructions (b). After that, subjects are asked to rate the images from three aspects (c), while carrying out tests (d) to control the annotation quality.
@@ 3.2 Design of the Human Evaluation
3.2.1 Choice of Naturalness-related Factors.
We conduct the naturalness evaluation from two perspectives, i.e., the low-level technical perspective and the high-level rationality perspective, as follows.
@@
(T-1) Luminance: Unrecognizable regions due to extremely high/low brightness.
@@
Participants and Apparatus. To ensure the comprehensiveness and reliability of the evaluation, we recruited 30 participants (18 male, 12 female, age=22.6±3.1) from campus, all with normal (corrected) eyesight. We conduct the subjective studies in-lab to ensure that all subjects have a clear and consistent understanding of all factors. Each participant is compensated \$240 for evaluating 6,049 images. All images are displayed on a 27-inch screen with a resolution of 2560$\times$1440 and a viewing distance of about 70cm. Note that we have addressed the ethical challenges involved in constructing such a database, by obtaining from each subject depicted in the database a signed and informed agreement, making it equipped with such legal and ethical characteristics.
Rating Strategy and Wording. We discuss the concrete form for human evaluation as follows. 1) Task-oriented absolute choice. Since the wording of questions and labels can significantly affect annotators’ labeling behavior, we abandon the traditional 3-point or 5-point Likert scale that only provides endpoint labels from worst to best and is too vague to describe the degree of naturalness. As a solution, we elaborate the questions and labels for three different perspectives to reduce subjectivity rather than using general ones (e.g. bad, good, or very good). 2) Pick up the main factor. Most existing subjective studies merely focus on the assessment of the overall score but neglect to explore the underlying factors. Therefore, we ask subjects to choose a primary factor that affects most for each perspective after rating the general scores, which enables us to investigate the correlation between each dimension and image naturalness.
Training, Testing, and Annotation. The workflow of the human evaluation is illustrated in Fig. 3. Before conducting the formal study, we manually generated 10 exemplars beyond the AGIN database for each dimension as training images to familiarize subjects with the goal of this evaluation. Subsequently, we instruct the subjects to rate the technical quality, rationality, and overall naturalness of each image from $\left\{1,2,3,4,5\right\}$, and select the main factors that affect the technical quality and rationality most. For testing, we randomly insert 10 golden images into each session as an inspection to ensure the quality of annotation. We defer more details of human evaluation and quality control to supplementary materials.
Figure 4: Data properties of AGIN. (a) The correlations between technical and rationality perspectives, (b) distributions of overall naturalness scores, and (c) the tendency of main factors chosen by participants across different ranges of naturalness scores.
Figure 5: More comparisons of image generation models in terms of naturalness-related factors. Zoom-in for better visualization.
@@ 3.3 Insights
Table 2: Correlation between different perspectives and overall naturalness in AGIN.
@@
$\mathrm{MOS_{T}+MOS_{R}}$ $\mathrm{0.145MOS_{T}+0.769MOS_{R}}$
@@
We first provide a visualization of the data properties in AGIN database (Fig. 4a), from which we can observe the inner correlation between the technical and rationality perspectives. Tab. 2 lists the quantitative results of Spearman and Pearson correlation between different perspectives, where the mean technical score, mean rationality score, and mean naturalness score are denoted as $\mathrm{MOS_{T}}$, $\mathrm{MOS_{R}}$, and $\mathrm{MOS}$, respectively. It can be observed that the two perspectives affect naturalness unequally (i.e., rationality has a greater impact on the overall naturalness than technical perspective, while the linear weighted sum of two perspectives can better approximate the overall naturalness than any single form). This could lead to biased naturalness evaluation unwittingly when using mainstream IQA models that follow the overall MOS regression strategy.
Insight ❷: Factors in two perspectives are related, but have disparate impacts on the overall naturalness.
Fig. 4c shows the proportion of each factor in different ranges of naturalness scores, where T-Null and R-Null indicate that there exist no factors affecting naturalness or the subjects have difficulty in choosing the main factors. First, T-Null and R-Null are more prevalent in images with preferable naturalness ($\mathrm{MOS}\in[4,5]$), indicating the impact of technical and rationality perspectives on naturalness. Additionally, we notice that humans are more sensitive to generated artifacts (T-5) and blur (T-4) in terms of technical quality while focusing more on the existence (R-1) of the image contents in terms of rationality. Furthermore, we find especial high proportions of artifacts (T-5), existence (R-1), and layout (R-3) in the case of poor naturalness ($\mathrm{MOS}\in[1,2]$), suggesting that they are important naturalness factors for AGIs to take into account. Considering the interrelation between technical quality and rationality, we speculate that severe artifact distortions may lead to irrational content and chaotic layout. For each generative task in AGIN, we calculate the occurrence frequency of each factor, as shown in Fig. 5. We also provide examples with varying degrees of effect for each dimension in Fig. 6, to better illustrate the manifestations of the naturalness problem in different dimensions. Overall, these newly contributed dimensions describe the naturalness concerns of AGIs, some of which have never been encountered in the conventional IQA domain, providing reliable intuitions for developing objective naturalness assessment models.
Figure 7: Framework of the proposed JOINT and JOINT++. It consists of the technical prior branch (Sec. 4.1) and the rationality perceiving branch (Sec. 4.2) with indirect and fine-grained supervisions strategies (Sec. 4.3).